OpenStudy (exploradora1011):

Solve the following equation algebraically: X^2=20

OpenStudy (daniellelovee):

what is the opposite of an exponent to the second value

OpenStudy (exploradora1011):

square?

OpenStudy (exploradora1011):

root

OpenStudy (queenie):

taking the sqrt of both sides will eliminate the ^2 on he left side

OpenStudy (daniellelovee):

correct

OpenStudy (exploradora1011):

These are my answer options though a. 4.47 b. -3.97, 3.97 c. -10, 10 d. -4.47, 4.47

OpenStudy (daniellelovee):

yes because you don't know if x is negative or not so you have two possible answers :)

OpenStudy (daniellelovee):

did you plug the sqrt of 20 on the calculator already?

OpenStudy (exploradora1011):

Yes I did, Its D

OpenStudy (daniellelovee):

correct :)

OpenStudy (queenie):

yep

OpenStudy (exploradora1011):

Thank you both so much!

OpenStudy (mathstudent55):

Below, x is the variable, and k is a number. The solution of the equation \(x^2 = k\) is \(x = \pm \sqrt k\)

OpenStudy (mathstudent55):

With the above in mind, \(x^2 = 20\) \(x = \pm \sqrt 20\)